OpenStudy (anonymous):

rationalize the denominator and simplify 3+√11/4-√11

OpenStudy (anonymous):

\[(3+\sqrt{11})/(4-\sqrt{11}) \] To rationalize denominator:\[(3+\sqrt{11})/(4-\sqrt{11}) * (4+\sqrt{11})/(4+\sqrt{11})\]\[(23+7\sqrt{11})/5\]